OpenStudy (summersnow8):

What would be the weight of a 59.1-kg astronaut on a planet with the same density as Earth and having twice Earth's radius? I am getting 290, but I dont know if that is correct.. A) 580N B) 290N C) 1160N D) 2320N E) 1200N

OpenStudy (summersnow8):

hat would be the weight of a 59.1-kg astronaut on a planet twice as massive as Earth and having twice Earth's radius? 290 What would be the weight of a 59.1-kg astronaut on a planet with the same density as Earth and having twice Earth's radius? 1160 I did \[\frac{ (6.67 x 10^{-11})(5.98 x 10^{24})(59.1) }{ ((2)(6.37 x 10^{2})) } = 289.56\]

OpenStudy (fifciol):

In other words weight is the gravitational force that acts on you \[F=\frac{ GMm }{ r^2 }\] where r is the distance between the mass m and center of mass M. In our case it's simply R of the Earth Mass of the Earth coan be written as \[M=\rho V=\rho \frac{ 4\pi R^3 }{3 }\] V is the volume of Earth(sphere) so we can write down equation of force in new form: \[F=\frac{ G4\pi R^3m }{ 3R^2 }=\frac{ G4\pi Rm }{ 3 }\] so this is a weight of astronaut on Earth ( proportional to radius of planet) On planet with twice Earth's radius we have twice this value: \[F=2*\frac{ G4\pi R^3m }{ 3R^2 }\] so astronaut would weigh twice than on Earth weigh of astronaut on Earth is 59,1* 9,81=579,771N weigh of astronaut on planet with twice Earth's radius is 2*59,1*9,81=1159,542 N

OpenStudy (summersnow8):

oh okay! thanks so much!!!

OpenStudy (fifciol):

yw:)